OpenStudy (christos):

Find the period of a function f(x) = 2cos(4x + 3Π)

hero (hero):

The period of a periodic function is the length it takes for the function to repeat itself again.

OpenStudy (christos):

can you please tell me how to find a period? Or an example of any sort? That would really help

hero (hero):

\[P = \frac{2\pi}{B}\]

OpenStudy (christos):

what is B?

hero (hero):

B = no. of cycles from 0 to \(2 \pi\)

hero (hero):

How many cycles are there?

OpenStudy (christos):

B= no ? no as what? And I dont know how many cycles there are, every information I got I posted it here

hero (hero):

There's two ways to tell. 1. From the equation: Acos(Bx + C) + D 2. From a graph: https://www.desmos.com/calculator/giaibd1bfz

hero (hero):

Did you figure out the period?